OpenStudy (anonymous):

what equsion is related to this equasion w/9=14 a.w=14/9 b.w=19*9 c.w=14-9 d.w=14+9

OpenStudy (anonymous):

If the given equation is\[\frac{ w }{ 9 } = 14\]then none of the answers are correct. To isolate \(w\) (solve for \(w\)), you would multiply both side of the equation by 9.

OpenStudy (anonymous):

Perhaps your answer b has a typo?

OpenStudy (anonymous):

acttuly the slach mean devide

OpenStudy (anonymous):

I understand but my comments still stand.

OpenStudy (anonymous):

ok

OpenStudy (welshfella):

look like there's a typo in option B

OpenStudy (anonymous):

As stated above.

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

Then there is no correct answer given.

OpenStudy (anonymous):

how

OpenStudy (welshfella):

oh sorry - I didnt notice your post @ospreytripl

OpenStudy (anonymous):

its ok

OpenStudy (anonymous):

Take \[\frac{ w }{ 9 } = 14\]and multiply both sides by 9. Can you do that?

OpenStudy (anonymous):

no math is hard for me

OpenStudy (anonymous):

Multiplying both side by 9 looks like\[\frac{ w }{ 9 } \times 9 = 14 \times 9\]Now on the left hand side the 9's will cancel, since \(\frac{9}{9}=1\), leaving\[w = 14 \times 9\]which is not one of the given answers. Understand?

OpenStudy (anonymous):

I still think there must be a typo in option b

OpenStudy (welshfella):

yes i agree

OpenStudy (anonymous):

oh wait its sopposed to be 14 not 19